OpenStudy (anonymous):

perform operation and write answer in lowest term 7z square +49z+70 / 16z square+ 72z-40 divided by 3z+6/4z square-1

OpenStudy (campbell_st):

factor the quadratics flip and mutiply the denominator \[\frac{7(z^2 + 7z + 10)}{8(2z^2 + 9z - 5)} \times \frac{4z^2 - 1}{3(z + 2)}\] hope this helps